Circular No. 7265

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 S3 (LINEAR) Additional astrometry is published on MPEC 1999-S18, from whence

the following preliminary parabolic orbital elements are taken:

T = 1999 Nov. 6.073 TT  ω = 42.^{o}810
Ω = 12.476 }2000.0
q = 1.92488 AU  i = 71.679
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Sept. 19 1^{h}41.^{m}16 +42^{o}23.^{′}8 1.229 2.014 128.^{o}6 23.^{o}0 13.5
24 1 23.34 +46 19.7 1.194 1.997 130.5 22.4 13.4
29 1 01.17 +50 01.1 1.172 1.982 131.3 22.3 13.3
Oct. 4 0 34.49 +53 15.2 1.162 1.968 130.8 22.6 13.3
9 0 03.81 +55 50.1 1.166 1.956 129.1 23.4 13.2

GM SAGITTARII AND SAX J1819.3–2525 = XTE J1819–254 Further to their report on IAUC 7254, R. M. Hjellming et al. write:

“Continued VLA observations of GM Sgr show that, since it became optically thin on Sept. 17.0 UT, the radio fluxes have decayed with a power law and spectrum approximated by 0.33(JD–2451437)^{−2.05}f^{−0.66}, where f is the frequency in GHz, reaching 3.9, 1.8, and 1.1 mJy at 1.49, 4.9, and 8.5 GHz on Sept. 22.0. VLA observations at 21 cm on Sept. 17.0 show strong H I absorption, identical within the noise to that seen against a nearby extragalactic calibrator (J1820–2528), suggesting that GM Sgr is at least a few hundred parsecs away; these data give no constraints on the maximum distance. The corresponding total H I column density is 1.8 × 10^{21} (T_{s}/100 K) cm^{−2}, where T_{s} is the spin temperature. On Sept. 16.02 the 0.41-Jy source at 4.9 GHz was resolved into a 0.34-Jy source (at the position reported on IAUC 7254) and extensions roughly 0′′.25 north and south. On Sept. 17.93, a 14.9-GHz image obtained with ‘fast-phase’ referencing to J1820– 2528 showed a ∼ 4-mJy source at the IAUC 7254 radio position and a ∼ 9-mJy source located ∼ 0′′.25 south, at the same location as the southern extension on Sept. 16.02. Observations on Sept. 22.00 and 24.1 show only this southern source. Accurate optical astrometry is needed to show where GM Sgr lies with respect to the radio emission. This would determine whether the Sept. 22–24 radio source should be interpreted as unusually long-lived ejecta, perhaps interacting with some surrounding material (cf. XTE J1748–288), or simply residual emission from the central object.”

1999 September 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green